[Hallc_running] [New Logentry] Hall C: status update - hall locking up tonight

gaskelld at jlab.org gaskelld at jlab.org
Mon Dec 4 09:20:03 EST 2017

Logentry Text:
Since yesterday, the program has changed a little - MCC will attempt to setup 1 pass beam to Hall C tonight (they have already delivered beam to the 3C05 dumplette). That means shift crews for Monday swing and Tuesday owl (which are "on call") should plan to show up for shifts.

We are also planning to condense the hydrogen target later today (the target group is coordinating with cryo to make sure this is ok). This allow some flexibility for Hall A since we cannot cool both Hall A and Hall C at the same time, and some time for recovery before cooldowns is desired.

Andy Kenyon is working on the SHMS collimator motion right now.  If it cannot be repaired today, we will open up again tomorrow to work on it some more.

If things go well with beam setup tonight, we may start Jay Benesch's "beam steering" test plan, so SHMS remote rotation and magnet setting will be needed.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3499353
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171204/5b91328a/attachment.html>

More information about the Hallc_running mailing list